OpenStudy (eamier):

a>0. integrate from 0 to infinity of e^(-ax^2)

OpenStudy (anonymous):

Uh, I think it's \(\large \sqrt{\frac{\pi}{4a}}\).

OpenStudy (eamier):

i cannot understand it, how we know that

OpenStudy (anonymous):

I thought you wouldn't. Do you know \(\large \int_0^{\infty}e^{-x^2}\mathrm{d}x\) ?

OpenStudy (eamier):

no.

OpenStudy (anonymous):

Ok, so I will show you how to find that, and then you can find your answer easily.

OpenStudy (experimentx):

should have put 2 outside the root

OpenStudy (anonymous):

So consider the following integral:\[\int_0^{\infty}\int_0^{\infty} e^{-x^2-y^2}dy~dx=\int_0^{\infty}\int_0^{\infty} e^{-x^2}e^{-y^2}dy~dx=\int_0^{\infty}e^{-x^2}\int_0^{\infty} e^{-y^2}dy~dx\]\[=\left (\int_0^{\infty}e^{-x^2}dx\right ) \left (\int_0^{\infty} e^{-y^2}dy\right ) =\left (\int_0^{\infty}e^{-x^2}dx\right )^2\]

OpenStudy (eamier):

in second last step, do we consider y = x?

OpenStudy (anonymous):

@eamier Well, x and y are both "dummy variables". The value of a definite integral does NOT depend on what variable we are using. It is simply a number. It depends on the function and the region of integration. So, if we find the volume under the surface e^(-x^2-y^2) in the first quadrant, we have found the square of our desired integral. (I hope you understand volume integrals. If not, look in your textbook or whatevs.) Now \(-x^2-y^2=-\left ( x^2+y^2\right ) =-\left ( \sqrt{x^2+y^2}\right ) ^2=-r^2\), where r is the distance from the origin, used in polar coordinates.

OpenStudy (anonymous):

So the surface depends only on r. Not on Θ. You should realize that this means the surface is a surface of revolution. You probably studied such surfaces in Calc 1. What we have is a function z=f(r) rotated around the z axis.

OpenStudy (anonymous):

But we only need one fourth of the entire volume.

OpenStudy (anonymous):

Do you think you can get it from there?

OpenStudy (eamier):

this really has nothing to do with when you integrate e, divide by the derivative of it power..i really cant imagine it. but how we get the pi

OpenStudy (anonymous):

Or take: \[\int_0^{\pi/2}\int_0^{\infty}e^{-r^2}r~dr~d\theta\]This is a simple polar integral which represents the volume under \(e^{-x^2-y^2}\) in the first quadrant. I know it may seem strange, but the math is flawless. You get the pi because multiplying \(e^{-x^2}\) with \(e^{-y^2}\) just so happens to give you \(e^{-r^2}\), and anything to do with polar coordinates usually has to do with pi. You will find that the above integral is equal to \(\frac{\pi}{4}\). Once you know that integral, you can use simple substitutions to find your original integral. Do you need help with the above integral?

OpenStudy (eamier):

i tried but cant do even this part\[\int\limits_{0}^{\infty}e ^{-r ^{2}}rdr\].

OpenStudy (experimentx):

let r^2 = u 2rdr = du

OpenStudy (eamier):

done. tq